The length of a rectangle is 3in. more than twice its width. If the perimeter of the rectange is 30in., find the dimensions of the rectangle. Choose the answer with the correct equations used to solve this problem.

Answer by nerdybill(7384)

Let w = width of rectangle
then
2w+3 = length of rectangle
.
2(w + 2w+3) = 30
2(3w+3) = 30
6w+6 = 30
w+1 = 5
w = 4 inches (width)
.
length:
2w+3 = 2(4)+3 = 8+3 = 11 inches (length)